### v2

@@ Abstract
* Equal contribution;   🖂 Corresponding author (wangwenhai362@gmail.com)
Transformers have revolutionized computer vision and natural language processing, but their high computational complexity limits their application in high-resolution image processing and long-context analysis. This paper introduces Vision-RWKV (VRWKV), a model adap-ted from the RWKV model used in the NLP field with necessary modifications for vision tasks. Similar to the Vision Transformer (ViT), our model is designed to efficiently handle sparse inputs and demonstrate robust global processing capabilities, while also scaling up effectively, accommodating both large-scale parameters and extensive datasets. Its distinctive advantage lies in its reduced spatial aggregation complexity, which renders it exceptionally adept at processing high-resolution images seamlessly, eliminating the necessity for windowing operations. Our evaluations demonstrate that VRWKV surpasses ViT’s performance in image classification and has significantly faster speeds and lower memory usage processing high-resolution inputs. In dense prediction tasks, it outperforms window-based models, maintaining comparable speeds. These results highlight VRWKV’s potential as a more efficient alternative for visual perception tasks. Code is released at https://github.com/OpenGVLab/Vision-RWKV.
Keywords:
RWKV Visual Perception Linear Attention
@@ 1 Introduction
Vision Transformers (ViTs) [12, 49, 52, 44, 18], renowned for their flexibility and global information processing capabilities, have established new benchmarks in a variety of vision tasks in the past few years. However, the quadratic computational complexity associated with ViTs limits their ability to efficiently process high-resolution images and lengthy sequences, posing a significant barrier to their broader application. As a result, the exploration of a vision architecture that integrates the versatility and comprehensive processing strengths of ViTs, while reducing computational demands, has emerged as a crucial area of research.
Figure 1: Performance and efficiency comparison of Vision-RWKV (VRWKV) and ViT. (a) $\rm AP^{b}$ comparison of VRWKV and ViT [49] with window attention and global attention on the COCO [27] dataset. (b) Inference speed comparison of VRWKV-T and ViT-T across input resolutions ranging from 224 to 2048. (c) GPU memory comparison of VRWKV-T and ViT-T across input resolutions from 224 to 2048.
@@
In recent developments within natural language processing (NLP), models like RWKV [35] and Mamba [15] have emerged as popular solutions for achieving heightened efficiency and processing lengthy texts. These innovative models have demonstrated attributes similar to transformers [9, 40, 43, 28, 38, 39, 3, 25, 45] in NLP tasks, including the ability to handle long-range dependencies and parallel processing. Furthermore, they have also proven to be scalable, performing well with large-scale NLP datasets. Considering the significant differences between image and text modalities, it remains challenging to envision these methods entirely supplanting ViTs for vision-related tasks. It is imperative to conduct an in-depth analysis of how these models are applied to vision tasks, examining their scalability concerning data and parameters, their efficiency in handling sparse visual data, such as masked image modeling, and the necessary techniques to ensure model stability during scaling up.
@@
In this work, we introduce Vision-RWKV (VRWKV), which is designed to adapt the RWKV architecture for vision tasks. This adaptation preserves the core structure and benefits of RWKV [35] while integrating critical modifications to tailor it for processing visual data. Specifically, we introduce a quad-directional shift (Q-Shift) tailed for vision tasks and modify the original causal RWKV attention mechanism to a bidirectional global attention mechanism. The Q-Shift operation expands the semantic range of individual tokens, while the bidirectional attention enables the calculation of global attention within linear computational complexity in an RNN form forward and backward. We primarily make modifications to the exponent in the RWKV attention mechanism, releasing the limitations of the decay vector and transforming the absolute positional bias into a relative bias. These changes enhance the model’s capability while ensuring scalability and stability. In this way, our model inherits the efficiency of RWKV in handling global information and sparse inputs, while also being able to model the local concept of vision tasks. We implement layer scale [50] and layer normalization [2] where needed to stabilize the model’s output across different scales. These adjustments significantly improve the model’s stability when scaling up to a larger size.
@@
Building on the aforementioned design, we develop a range of VRWKV models with different model scales, spanning from the VRWKV-Tiny (6M) to the VRWKV-Large (335M). These models are trained using large-scale datasets such as ImageNet-1K [8] and ImageNet-22K [8]. We train them using both common supervised classification and sparse input MAE methods [18] and evaluate their performance on visual perception tasks, including classification, detection, and segmentation. Under the same settings, VRWKV has comparable performance to ViT in these tasks with lower computational costs while maintaining stable scalability. This achievement enables VRWKV training parallelism, high flexibility, excellent performance, and low inference cost simultaneously, making it a promising alternative to ViT in a wide range of vision tasks, particularly in high-resolution scenarios.
@@
(1) We propose VRWKV as a low-cost alternative to ViT, achieving comprehensive substitution with lower computational costs. Our model not only retains the advantages of ViT, including the capability to capture long-range dependencies and flexibility in handling sparse inputs but also reduces complexity to a linear level. This significant reduction eliminates the need for window-based attention in processing high-resolution images, making VRWKV a more efficient and scalable solution for vision tasks.
(2) To adapt to vision tasks, we introduce bidirectional global attention and a novel token shift method called Q-Shift, enabling the achievement of linear complexity in global attention. To ensure stable scalability, we make several efforts, including using a relative positional bias in the attention mechanism to avoid overflow, adopting layer scale in our model, and adding extra layer normalization in the calculation of key matrices.
(3) Our model surpasses window-based ViTs and is comparable to global attention ViTs, demonstrating lower FLOPs and faster processing speeds as resolution increases. Notably, VRWKV-T achieves 75.1% top-1 accuracy trained only on the ImageNet-1K [8], outperforming DeiT-T [49] by 2.9 points. With large-scale parameters (i.e., 335M) and training data (i.e., ImageNet-22K), the top-1 accuracy of VRWKV-L is further boosted to 86.0%, which is higher than ViT-L [12] (86.04 vs. missing 85.15). In addition, on COCO [27], a challenging downstream benchmark, our best model VRWKV-L achieves 50.6% box mAP, 1.9 points better than ViT-L (50.6 vs. missing 48.7).
@@ 2.1 Vision Encoder
Recent advances in vision encoders have significantly pushed the boundaries of computer vision, demonstrating remarkable performance across a range of tasks. Convolutional neural networks (CNNs) served as the foundational model in computer vision. The advancement of computational resources, such as GPUs, has enabled the successful training of stacked convolutional blocks like AlexNet [23] and VGG [41] on large-scale image classification datasets (e.g., ImageNet [8]). This development paved the way for deeper and more sophisticated convolutional neural architectures, including GoogleNet [48], ResNet [20], and DenseNet [22].
In addition to these innovations, significant advancements have also been made with architectures like SENet [21], which introduced a channel attention mechanism to enhance model sensitivity to informative features. Similarly, SKNet [26] merged multiple kernel sizes to adjust the receptive field adaptively. Further extending the CNN paradigm, recent models such as RepLKNet [11] and ConvNeXt [31] have refined the convolutional layers to improve efficiency and accuracy, while InternImage [55] explored the strategies to scale up the convolution-based vision model.
Drawing inspiration from the effectiveness of self-attention layers and transformer architectures in the NLP field, the Vision Transformer (ViT) [12] applied a transformer framework on image patches, offering a global receptive field and dynamic spatial aggregation. Due to the quadratically increasing computational complexity of the vanilla attention mechanism, approaches like PVT [56, 57] and Linformer [54] implemented global attention on down-sampled feature maps, whereas other approaches like Swin [58] and HaloNet [51, 6] introduced sampling techniques to enlarge the receptive field.
Another research direction involved replacing self-attention layers in models with linear complexity layers. Representative works in this domain include LongNet [10], RWKV [35], RetNet [47], and Mamba [15], though few have concentrated on vision-related applications. Concurrently, attempts like Vim [63] and VMamba [29] have sought to integrate these linear attention layers into the vision domain. However, these endeavors have only been experimented with on small-scale models, and it remains uncertain whether their efficiency can scale up to larger models.
@@ 2.2 Feature Aggregation Mechanism
The research on feature aggregation has received significant attention in the field of artificial intelligence. For visual data processing, convolutional operators [24], known for their parameter sharing and local perception, enabled efficient handling of large-scale data through sliding computation. Despite their advantages, traditional CNN operators faced challenges in modeling long-range dependencies. To overcome this issue, advanced convolutional operators, such as the deformable convolution [5, 64, 60], have improved the flexibility of CNN operators, enhancing their long-range modeling capability.
@@
As for the field of NLP, RNN-based operators [13, 34, 37] have historically dominated because of their effectiveness in sequence modeling. RNNs and LSTMs excel in capturing temporal dependencies, making them suitable for tasks requiring an understanding of sequence dynamics. Subsequently, a significant shift occurred. The introduction of the transformer architecture [52] marked a turning point, with both NLP and computer vision fields shifting focus toward attention-based feature aggregation. The global attention mechanism overcomes the limitations of CNNs in modeling long-range dependencies and the shortcomings of RNNs in parallel computation while coming at a high computational cost.
@@
To address the high computational cost of attention operators while modeling long sequences, researchers have introduced innovations such as window attention and spatial reduction attention. Window attention [30, 51, 6] restricts the self-attention computation within local windows, drastically reducing the computational complexity while preserving the receptive field through window-level interaction. Spatial reduction attention [56, 57], on the other hand, reduces the dimensionality of the feature space before applying the attention mechanism, effectively decreasing the computational requirements without significantly degrading the model’s performance.
@@
In addition to the efforts to optimize the global attention mechanism, various operators with linear complexity [35, 47, 15, 36, 46] have also been explored. For instance, RWKV [35] and RetNet [47] employed exponential decay to model global information efficiently while SSMs [16, 17, 42, 53] also exhibited linear complexity concerning sequence length and modification in Mamba [15] enable them to be input-dependent. Besides, XCA [1] achieved linear complexity by calculating the cross-variance between input tokens. However, the low efficiency of information interaction between tokens makes the need for the assistance of additional modules necessary to complete a comprehensive feature aggregation. Despite some concurrent efforts [29, 63, 14], adapting these NLP-derived techniques to vision tasks remains a challenge in maintaining stable training across larger and more complex vision models.
@@
Figure 2: Overall architecture of Vision-RWKV (VRWKV). (a) The VRWKV architecture includes $L$ identical VRWKV encoder layers, an average pooling layer, and a linear prediction head. (b) The details of the VRWKV encoder layer. Q-Shift denotes the quad-directional shift method tailed for vision tasks. The “Bi-WKV” module served as a bidirectional RNN cell or a global attention mechanism.
@@ 3.1 Overall Architecture
In this section, we propose Vision-RWKV (VRWKV), an efficient vision encoder with a linear complexity attention mechanism. Our principle is to preserve the advantages of the original RWKV architecture [35], making only necessary modifications to enable its flexible application in vision tasks, supporting sparse input, and ensuring the stability of the training process after scaling up. An overview of our VRWKV is depicted in Fig. 2.
@@
In each layer, tokens are first fed into the spatial-mix module which plays the role of a global attention mechanism. Specifically, as shown in Fig. 2(b), the input tokens are first shifted and fed into three parallel linear layers to obtain the matrices $R_{s},K_{s},V_{s}\in\mathbb{R}^{T\times C}$:
@@
Here, $K_{\text{s}}$ and $V_{\text{s}}$ are passed to calculate the global attention result $wkv\in\mathbb{R}^{T\times C}$ by a linear complexity bidirectional attention mechanism and multiplied with $\sigma(R)$ which controls the output $O_{\text{s}}$ probability:
@@
Operator $\sigma$ denotes the sigmoid function, and $\odot$ means an element-wise multiplication is applied. The $\rm Q\mbox{-}Shift$ is a token shift function specially designed for the adaption of vision tasks. After an output linear projection, features are then stabilized using layer normalization [2].
@@
Here, $V_{\text{c}}$ is a linear projection of $K$ after the activation function and the output $O_{\text{c}}$ is also controlled by a gate mechanism $\sigma(R_{\text{c}})$ before the output projection:
@@
Simultaneously, residual connections [20] are established from the tokens to each normalization layer to ensure that training gradients do not vanish in deep networks.

### v3

@@ Abstract
Transformers have revolutionized computer vision and natural language processing, but their high computational complexity limits their application in high-resolution image processing and long-context analysis. This paper introduces Vision-RWKV (VRWKV), a model that builds upon the RWKV architecture from the NLP field with key modifications tailored specifically for vision tasks. Similar to the Vision Transformer (ViT), our model demonstrates robust global processing capabilities, efficiently handles sparse inputs like masked images, and can scale up to accommodate both large-scale parameters and extensive datasets. Its distinctive advantage is its reduced spatial aggregation complexity, enabling seamless processing of high-resolution images without the need for window operations. Our evaluations demonstrate that VRWKV surpasses ViT’s performance in image classification and has significantly faster speeds and lower memory usage processing high-resolution inputs. In dense prediction tasks, it outperforms window-based models, maintaining comparable speeds. These results highlight VRWKV’s potential as a more efficient alternative for visual perception tasks. Code and models are available at https://github.com/OpenGVLab/Vision-RWKV.
@@ 1 Introduction
Vision Transformers (ViTs) (Dosovitskiy et al., 2020; Touvron et al., 2021a; Vaswani et al., 2017; Steiner et al., 2021; He et al., 2021), renowned for their flexibility and global information processing capabilities, have established new benchmarks in a variety of vision tasks in the past few years. However, the quadratic computational complexity associated with ViTs limits their ability to efficiently process high-resolution images and lengthy sequences, posing a significant barrier to their broader application. As a result, the exploration of a visual architecture that integrates the versatility and comprehensive processing strengths of ViTs, while reducing computational demands, has emerged as a crucial area of research.
@@
In recent developments within natural language processing (NLP), models with linear feature aggregation (or called “linear attention”) mechanisms like RWKV (Peng et al., 2023) and Mamba (Gu & Dao, 2023) have emerged as popular solutions for achieving heightened efficiency and processing lengthy texts. These innovative models have demonstrated attributes similar to transformers (Devlin et al., 2018; Raffel et al., 2019; Smith et al., 2022b; Liu et al., 2019; Radford et al., 2018; 2019; Brown et al., 2020; Lewis et al., 2019) in NLP tasks, including the ability to handle long-range dependencies and parallel processing. Furthermore, they have also proven to be scalable, performing well with large-scale NLP datasets. Expanding these techniques into the visual domain shows promise in addressing the computational cost challenge encountered by ViTs.
To develop a vision model incorporating a linear attention mechanism based on the aforementioned methods, while ensuring high capacity for large-scale image data and diverse visual tasks, several critical issues need to be addressed. Firstly, the design of spatial feature aggregation operations needs to be reconsidered, taking into account the differences between image and text modalities. For instance, a redesign of kernels and rewriting at the CUDA level are necessary for attention mechanisms with a causal receptive field in models like RWKV. Furthermore, the issues of gradient vanishing or exploding tend to arise gradually as the model scales up, resulting in unstable training with large parameter sizes and large-scale datasets. For example, Vision Mamba (Zhu et al., 2024) only gave appropriate results on models with less than 30M parameters. It is important to conduct an in-depth study of how linear attention models can be applied to vision tasks effectively, including examining the scalability concerning data and parameters, assessing the efficiency in handling sparse visual data, and implementing necessary techniques to ensure model stability during scaling up.
@@
Based on these points, we introduce Vision-RWKV (VRWKV). Our approach preserves the core structure and benefits of RWKV (Peng et al., 2023) while incorporating essential changes to process visual data efficiently. Specifically, we design a quad-directional shift (Q-Shift) tailed for vision tasks and modify the original causal RWKV attention mechanism to a bidirectional global attention mechanism (Bi-WKV). The Q-Shift operation expands the semantic range of individual tokens, while the Bi-WKV enables the calculation of global attention within linear complexity in an RNN-form forward and backward. We primarily modify the exponent in the RWKV attention, releasing the limitations of the decay vector and transforming the absolute positional bias into a relative bias. These changes enhance the model’s capability while ensuring scalability and stability. In this way, our VRWKV inherits the efficiency of RWKV in handling global information and sparse inputs, while also being able to model the local concept of vision tasks. Additionally, due to severe instability encountered when scaling up the model, we explored a series of measures (Touvron et al., 2021b; Ba et al., 2016) to stabilize the model’s outputs. These adjustments significantly improve the model’s training stability when scaling up to a larger size.
Figure 1: Performance and efficiency comparison of Vision-RWKV (VRWKV) and ViT. (a) Bounding box average precision ($\rm AP^{b}$) comparison of VRWKV and ViT (Touvron et al., 2021a) with window attention and global attention on the COCO (Lin et al., 2014) dataset. (b) Inference speed comparison of VRWKV-T and ViT-T across input resolutions ranging from 224 to 2048. (c) GPU memory comparison of VRWKV-T and ViT-T across input resolutions from 224 to 2048.
@@
Building on the aforementioned design, we develop a range of VRWKV models with different model scales, spanning from the VRWKV-Tiny (6M) to the VRWKV-Large (335M). These models are trained using large-scale datasets such as ImageNet-1K (Deng et al., 2009) and ImageNet-22K (Deng et al., 2009). We train them using both common supervised classification and sparse input method MAE (He et al., 2021) and evaluate their performance on visual perception tasks, including classification, detection, and segmentation. Under the same settings, VRWKV has comparable performance to ViT in these tasks with lower computational costs while maintaining stable scalability. This achievement enables VRWKV training parallelism, high flexibility, excellent performance, and low inference cost simultaneously, making it a promising alternative to ViT in a wide range of vision tasks, particularly in high-resolution scenarios.
@@
(1) We propose VRWKV as a cost-effective alternative to ViT, offering a comprehensive substitute with lower computational requirements. Our model retains ViT’s strengths, such as capturing long-range dependencies and handling sparse inputs flexibly, while reducing complexity to a linear scale. This reduction eliminates the need for window operation when processing high-resolution images, making VRWKV a more efficient and scalable solution for vision tasks.
(2) To support vision tasks, we develop a bidirectional global attention mechanism combined with a novel token shift method, Q-Shift, to achieve linear complexity in global attention. Additionally, we implement a set of tailored strategies—integrating relative positional bias, layer scale, and extra layer normalization—to tackle overflow issues and ensure stable, scalable training.
(3) Our model surpasses window-based ViTs and is comparable to global attention ViTs, demonstrating lower FLOPs and GPU memory cost with faster processing speeds as resolution increases, as shown in Figure 1. Notably, VRWKV-T achieves 75.1% top-1 accuracy trained only on the ImageNet-1K (Deng et al., 2009), outperforming DeiT-T (Touvron et al., 2021a) by 2.9 points. With large-scale parameters (i.e., 335M) and training data (i.e., ImageNet-22K), the top-1 accuracy of VRWKV-L is further boosted to 86.0%, which is higher than ViT-L (Dosovitskiy et al., 2020) (86.04 vs 85.15). In addition, on COCO (Lin et al., 2014), a challenging downstream benchmark, our best model VRWKV-L achieves 50.6% box mAP, 1.9 points better than ViT-L (50.6 vs 48.7).
@@ 2.1 Vision Encoder
Recent advances in vision encoders have significantly pushed the boundaries of computer vision, demonstrating remarkable performance across a range of tasks. Convolutional neural networks (CNNs) served as the foundational model in computer vision. The advancement of computational resources, such as GPUs, has enabled the successful training of stacked convolutional blocks like AlexNet (Krizhevsky et al., 2012) and VGG (Simonyan & Zisserman, 2014) on large-scale image classification datasets (e.g., ImageNet (Deng et al., 2009)). This development paved the way for deeper and more sophisticated convolutional neural architectures, including GoogleNet (Szegedy et al., 2015), ResNet (He et al., 2016), and DenseNet (Huang et al., 2017).
In addition to these innovations, significant advancements have also been made with architectures like SENet (Hu et al., 2018), which introduced a channel attention mechanism to enhance model sensitivity to informative features. Similarly, SKNet (Li et al., 2019) merged multiple kernel sizes to adjust the receptive field adaptively. Further extending the CNN paradigm, recent models such as RepLKNet (Ding et al., 2022) and ConvNeXt (Liu et al., 2022) have refined the convolutional layers to improve efficiency and accuracy, while InternImage (Wang et al., 2023b) explored the strategies to scale up the convolution-based vision model.
Inspired by the success of self-attention layers and transformer architectures in the NLP field, the Vision Transformer (ViT) (Dosovitskiy et al., 2020) applied a transformer framework on image patches, offering a global receptive field and dynamic spatial aggregation. Due to the quadratically computational complexity of the vanilla attention mechanism, approaches like PVT (Wang et al., 2021; 2022) and Linformer (Wang et al., 2020) implemented global attention on down-sampled feature maps, whereas other approaches like Swin (Wu et al., 2022) and HaloNet (Vaswani et al., 2021; Dai et al., 2022) introduced sampling techniques to enlarge the receptive field. Mini-InternVL (Gao et al., 2024) reduces the parameter size of ViT by employing knowledge distillation from a larger ViT to a smaller one, thereby achieving efficiency in the visual encoder.
Another research direction involved replacing self-attention layers in models with linear complexity layers. Representative works include LongNet (Ding et al., 2023), RWKV (Peng et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2023), though few have concentrated on visual applications. Concurrently, attempts like Vim (Zhu et al., 2024) and VMamba (Liu et al., 2024) have sought to integrate these linear attention layers into the vision domain. However, these endeavors have only been experimented with on small-scale models (parameters $\rm<30M$ for Vim and $\rm<100M$ for VMamba), leaving it uncertain whether their effectiveness extends to larger models.
@@ 2.2 Feature Aggregation Mechanism
The research on feature aggregation has received significant attention. For visual data processing, convolutional operators (LeCun et al., 1995), known for their parameter sharing and local perception, enabled efficient handling of large-scale data through sliding computation. Despite their advantages, traditional CNN operators faced challenges in modeling long-range dependencies. To overcome this issue, advanced convolutional operators, such as the deformable convolution (Dai et al., 2017; Zhu et al., 2019; Xiong et al., 2024), have improved the flexibility of CNN operators, enhancing their long-range modeling capability.
@@
As for the field of NLP, RNN-based operators (Elman, 1990; Memory, 2010; Qin et al., 2024) have historically dominated because of their effectiveness in sequence modeling. RNNs and LSTMs excel in capturing temporal dependencies, making them suitable for tasks requiring an understanding of sequence dynamics. Subsequently, a significant shift occurred. The introduction of the transformer architecture (Vaswani et al., 2017) marked a turning point, with both NLP and computer vision fields shifting focus toward attention-based feature aggregation. The global attention mechanism overcomes the limitations of CNNs in modeling long-range dependencies and the shortcomings of RNNs in parallel computation while coming at a high computational cost.
@@
To address these issues, researchers have introduced innovations such as window attention and spatial reduction attention. Window attention (Liu et al., 2021; Vaswani et al., 2021; Dai et al., 2022) restricts the self-attention computation within local windows, drastically reducing the computational complexity while preserving the receptive field through window-level interaction. Spatial reduction attention (Wang et al., 2021; 2022), on the other hand, reduces the dimensionality of the feature space before applying the attention mechanism, effectively decreasing the computational requirements without significantly degrading the model’s performance.
@@
In addition to the efforts to optimize the global attention mechanism, various operators with linear complexity have also been explored. For instance, RWKV (Peng et al., 2023) and RetNet (Sun et al., 2023) employed exponential decay to model global information efficiently while SSMs (Gu et al., 2021a; b; Smith et al., 2022a; Wang et al., 2023a) also exhibited linear complexity concerning sequence length and modification in Mamba (Gu & Dao, 2023) enable them to be input-dependent. Besides, XCA (Ali et al., 2021) achieved linear complexity by calculating the cross-variance between input tokens. However, the low efficiency of information interaction between tokens makes the need for the assistance of additional modules necessary to complete a comprehensive feature aggregation. Despite some concurrent efforts (Liu et al., 2024; Zhu et al., 2024; Fan et al., 2023), adapting these NLP-derived techniques to vision tasks remains a challenge in maintaining stable training across larger and more complex vision models.
@@
Figure 2: Overall architecture of VRWKV. (a) The VRWKV architecture includes $L$ identical VRWKV encoder layers, an average pooling layer, and a linear prediction head. (b) The details of the VRWKV encoder layer. Q-Shift denotes the quad-directional shift method tailed for vision tasks. The “Bi-WKV” module served as a bidirectional RNN cell or a global attention mechanism.
@@ 3.1 Overall Architecture
In this section, we propose Vision-RWKV (VRWKV), an efficient vision encoder with a linear complexity attention mechanism. Our principle is to preserve the advantages of the original RWKV architecture (Peng et al., 2023), making necessary modifications to enable its flexible application in vision tasks, supporting sparse input, and ensuring the stability of the training process after scaling up. An overview of our VRWKV is depicted in Figure 2.
@@
In each layer, tokens are first fed into the spatial-mix module which plays the role of a global attention mechanism. Specifically, as shown in Figure 2(b), the input tokens are first shifted and fed into three parallel linear layers to obtain the matrices $R_{s},K_{s},V_{s}\in\mathbb{R}^{T\times C}$:
@@
Here, the $\rm Q\mbox{-}Shift$ operator is a token shift function specially designed for the information exchange through nearby tokens according to the visual prior. $K_{\text{s}}$ and $V_{\text{s}}$ are then passed to calculate the global attention result, $wkv\in\mathbb{R}^{T\times C}$, by a linear complexity bidirectional attention mechanism, Bi-WKV, and multiplied with $\sigma(R)$ which controls the output $O_{\text{s}}$ probability:
@@
Operator $\sigma$ denotes the sigmoid function, and $\odot$ represents element-wise multiplication. The output features are then stabilized using layer normalization (Ba et al., 2016) following the linear projection.
@@
In the channel-mix module, $V_{\text{c}}$ is a linear projection of $K_{\text{c}}$ after the activation function and controlled by a gate mechanism $\sigma(R_{\text{c}})$. The output $O_{\text{c}}$ is the linear projection of the aforementioned result:
@@
Simultaneously, residual connections (He et al., 2016) are established from the tokens to each normalization layer to ensure that training gradients do not vanish in deep networks.